OpenStudy (anonymous):

Determine whether the sequence could be arithmetic. If so, find the common difference and the next term. 11/2, 11/3, 11/4, 11/5, 11/6 fractions

OpenStudy (anonymous):

now 2/11,3/11 etc are in AP so this could be in HP

OpenStudy (dumbcow):

look at the differences 11/3 - 11/2 = 22/6 - 33/6 = -11/6 11/4 - 11/3 = 33/12 - 44/12 = -11/12 they are different so this is Not an Arithmetic sequence...its not adding the same amount each time

OpenStudy (dumbcow):

however...you could say that the denominators form an arithmetic sequence, adding 1 each time